OpenStudy (anonymous):

algrebra question:

OpenStudy (misty1212):

ooh good!!!

OpenStudy (anonymous):

a common formula used to find the area of a triangle is A = bh/2, where b is the measurement of length of the base and h is the measurement of height. the area of a triangle is 15cm to the power of 2, and the base is 3 cm. use algebraic operations to solve an equation to find the measurement of the height of the triangle.

OpenStudy (misty1212):

wowee what a lot of words divide

OpenStudy (anonymous):

Ikr!

OpenStudy (misty1212):

\[A=\frac{bh}{2}\\ 15=\frac{3\times h}{2}\] lets see if we can get \(h\)

OpenStudy (michele_laino):

Please note that you have to solve this equation: A = (b*h)/2 in order to find h

OpenStudy (misty1212):

multiply 15 by the "reciprocal' of \(\frac{3}{2}\) reciprocal means flip it, so \[\frac{2}{3}\times 15=h\]

OpenStudy (anonymous):

is it 10/9 = h

OpenStudy (misty1212):

hmm no

OpenStudy (anonymous):

just 10 = h

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

two times fifteen is thirty, thirty divide by three is ten fifteen divided by three is five, five times two is ten

OpenStudy (misty1212):

either way

OpenStudy (anonymous):

got it!! so thats the answer?

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Thanks sooooo much! Mwaaaaaaaah :-*

OpenStudy (misty1212):

\[\huge \color\magenta\heartsuit\]